Some text to provide context.

Figure 1: 8 × 8, unit=0.8

more text, and picture with same scale, half as many vertical spaces Figure 2: 8 × 4, unit=0.8 same picture with half scale, full number vertical spaces

Figure 3: 8 × 8, unit=0.4

more text, and picture with full scale, half as many vertical spaces Figure 4: 8 × 4, unit=0.8 even more text, again the original picture,

Figure 5: 8 × 8, unit=0.8

and even more text, followed by picture same number vertical spaces but half the unit size Figure 6: 8 × 8, unit=0.4

4